News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 16 February 2001 11.00am EST

During the last week, Chandra completed the observing schedule as planned.

Chandra passed through the first 3 eclipses of the Spring 2001 season with nominal power and thermal performance. All three eclipses were classed as short eclipses (<40 minutes in duration). A software patch to configure for the upcoming medium eclipses will be uplinked on 2/17.

Real-time procedures completed last week included the insertion of the LETG on 2/14 and the re-dumping of SSR-A following a switch over to SSR-B after the recorder filled up due to station problems on 2 sequential passes. All data have been retrieved and the standard dump pattern has been restored. In addition, a check was performed of the CPE warm-start count that showed the value remained at 1.

The planned observations for the coming week are shown below and include observations of the Crab Nebula and NGC 3783 coordinated with the Hubble Space Telescope and two long observations of the Hubble Deep Field North totaling 250 ks.

------------------------------------------
Radiation Belt                       Feb 18 
Aql X-1                 ACIS-S       Feb 19 
HDF-N                   ACIS-I
Crab Nebula             ACIS-S       Feb 20 
Radiation Belt                       Feb 21 
HDF-N                   ACIS-I       Feb 21 
Radiation Belt                       Feb 24 
NGC 3783                ACIS-S/HETG
Radiation Belt                       Feb 26

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports